OpenStudy (anonymous):

16 2/3% of what number is 300

OpenStudy (anonymous):

I always use the method \[\frac{ Percent }{ 100 }=\frac{ is }{ of }\] therfore \[\frac{ 50/3 }{ 100 }=\frac{ 300 }{ x }\] then you cross multiply \[(50/3)x=3000\] isolate x and you have your answer

OpenStudy (anonymous):

i got 180 because i divide 3000 and (50/3)